# Time-Dependent Black Hole Lensing and the Unified Weak-to-Strong Deflection Framework

### v1

@@ Abstract
We present a fully analytical framework that unifies weak-field and strong-deflection lensing of light in a time-dependent, perturbed Schwarzschild spacetime. The spacetime dynamics are modeled by a single, axisymmetric, even-parity quasinormal mode with $\ell=2$, $m=0$ and complex frequency $\omega$. Working to first order in a small perturbation amplitude while keeping background null geodesics exact, we derive a time-dependent line-of-sight (Born) expression for the screen-plane deflection measured by a static observer at large radius. From the same integral, an asymptotic expansion yields the familiar weak-field $1/b$ law with a ringdown-frequency correction that drives a harmonic centroid wobble, whereas a near-photon-sphere expansion produces a time-dependent generalization of the logarithmic strong-deflection limit with modulated coefficients, including a small oscillation of the critical impact parameter. An observer tetrad built from the background static frame ensures that all screen-plane quantities—centroid motion, multi-image hierarchy and time delays, and photon-ring morphology—are gauge-safe at first order. We provide explicit matching across regimes, showing that the near-critical coefficients governing spacing and ring-radius modulations are encoded in the same Born kernel that controls the weak-field correction. The result is a coherent, purely theoretical account of how ringdown physics imprints on imaging observables without numerical ray tracing.
Gravitational lensing; quasinormal modes; Schwarzschild black hole; strong-deflection limit; photon ring.
@@ I Introduction
Gravitational lensing by compact objects remains a powerful probe of strong-field gravity and the nature of dark compact sources Zhong et al. (2025); Johnson et al. (2020); Kaiser and Jaffe (1997); Damour and Esposito-Farese (1998); Kopeikin and Schaefer (1999); Tsukamoto (2017); Vagnozzi et al. (2023); Allahyari et al. (2020); Khodadi et al. (2020); Molla et al. (2025); Ditta et al. (2025); Turakhonov et al. (2024); Alloqulov et al. (2025). Foundational analyses of Schwarzschild lensing and its extensions to naked singularities established the basic phenomenology of multiple images, relativistic loops, and their observational signatures Virbhadra and Ellis (2000, 2002). Early work also highlighted how additional fields can imprint themselves on lensing observables Virbhadra et al. (1998) and connected these effects to broader questions such as Seifert’s conjecture Virbhadra (1999). The differential‐geometric underpinning of photon surfaces furnished a rigorous framework for locating unstable light rings that control both strong lensing and shadow formation Claudel et al. (2001). Building on this, detailed predictions for relativistic images, time delays, and magnification centroids were developed and refined Virbhadra (2009); Virbhadra and Keeton (2008). Recent progress has quantified image distortions in Schwarzschild lensing and uncovered conservation properties of distortion measures, providing robust diagnostics that are less sensitive to astrophysical systematics Virbhadra (2022, 2024a). In parallel, precise relations between the photon-sphere radius and black-hole shadow have been extended to include cosmological constant corrections, clarifying how large-scale curvature perturbs near-horizon optics Adler and Virbhadra (2022). These theoretical advances feed directly into constraints on the compactness of supermassive dark objects at galactic centers, sharpening tests of black-hole paradigms with high-resolution imaging and timing data Virbhadra (2024b). Realistic black hole lensing is rarely vacuum: ambient plasma and magnetic fields induce frequency-dependent deflection, reshape caustics, and shift photon rings and shadows Morozova et al. (2013); Schee et al. (2017); Turimov et al. (2019). Such chromatic and dispersive effects persist across rotating and regular spacetimes, complicating image interpretation Morozova et al. (2013); Schee et al. (2017). Beyond GR, braneworld geometries predict systematic deviations in strong-deflection and retrolensing observables, offering complementary tests Abdujabbarov et al. (2017).
@@
Gravitational lensing by black holes spans a wide dynamic range from weak deflections in the asymptotic region to the strong-field regime near the photon sphere Virbhadra and Ellis (2000); Perlick and Tsupko (2022). In parallel, the spacetime itself can be time-dependent, for instance, during black-hole ringdown when quasinormal modes (QNMs) drive metric perturbations with characteristic complex frequencies Berti et al. (2009); Kokkotas and Schmidt (1999). These two features, wide spatial scales and intrinsic temporal modulation, are often modeled with separate tools that obscure their continuity. Our objective is to provide a single, purely analytical framework that treats both regimes within one description and that projects directly onto the observer’s screen in a gauge-safe manner.
@@
We concentrate on light propagation in the Schwarzschild geometry of mass $M$, perturbed at first order by a single axisymmetric even-parity Zerilli mode with angular numbers $\ell=2$, $m=0$ and complex frequency $\omega=\omega_{R}+i\omega_{I}$ with $\omega_{I}<0$ Zerilli (1970); Moncrief (1974). This choice provides the simplest dynamical driver of ringdown that preserves axial symmetry, minimizes technical overhead, and still captures the essential physics of a decaying spacetime oscillation. Throughout, we work perturbatively in a small, dimensionless amplitude $\varepsilon$ and keep the unperturbed null geodesics exact, so that all time dependence arises from the interaction of the QNM with the photon’s background trajectory Kopeikin and Schaefer (1999); Lewis and Challinor (2006).
@@
Our central construct is a time-dependent line-of-sight map that gives the deflection measured by a static observer at a large areal radius. The map is evaluated along the unperturbed null geodesic labeled by its impact parameter $b$. In the far field, it reproduces the familiar weak-bending law ($1/b$) Einstein (2005); Weinberg (1972); Will (2014) with an $\varepsilon$-suppressed, frequency-resolved correction that produces a harmonic motion of the image centroid on the screen. In the near-critical regime $b\to b_{c}^{+}$, with $b_{c}$ the Schwarzschild critical impact parameter, the same map yields a time-dependent generalization of the logarithmic strong-deflection relation Bozza (2002); Tsukamoto (2017). The coefficients of this relation become slowly varying functions of the observer’s time, all oscillating at the QNM frequency. The critical scale itself acquires a small modulation that controls the most singular response near the photon ring Chandrasekhar (1985); Johnson et al. (2020); Gralla et al. (2020); Akiyama et al. (2019, 2022).
@@
A key structural feature of our approach is the use of an observer tetrad fixed by the background static frame Moncrief (1974); Johnson et al. (2020). Screen angles are constructed directly from locally measured photon momenta. At first order in $\varepsilon$, this guarantees that centroid shifts, image positions, and ring morphology are insensitive to gauge transformations that would otherwise act on the metric perturbation. In this way, the imaging diagnostics we develop are bona fide observables at the order of interest.
@@
Recent work has begun to probe time-dependent gravitational lensing during black-hole relaxation. In particular, Zhong, Cardoso, and Chen Zhong et al. (2025) have used backward ray tracing in a relaxing Kerr background to show that the photon deflection inherits the characteristic ringdown pattern at intermediate times and exhibits a $t_{o}^{-3}$ late-time decay, thereby establishing the qualitative imprint of QNM dynamics on light propagation Price (1972). Earlier analyses of time-dependent deflection by gravitational waves or localized, retarded sources also demonstrated oscillatory signatures in lensing and timing, albeit in regimes far from the photon sphere and without a black-hole ringdown setting Kaiser and Jaffe (1997); Damour and Esposito-Farese (1998); Wang et al. (2021); Pantig and Övgün (2024); Khodadi et al. (2025). At the same time, the strong-deflection limit (SDL) for stationary spacetimes is well understood, including the logarithmic divergence and its coefficients ($\bar{a},\bar{b},b_{c}$) that control image hierarchies and photon-ring morphology Tsukamoto (2017). Classic studies of light propagation in nonstationary fields also delineate key limits: in particular, for localized sources, the wave-zone and intermediate-zone contributions cancel so that the leading time-dependent deflection is near-zone and falls as $b^{-3}$ (contrasting with the $1/b$ Born imprint we obtain from a global QNM on a black-hole background), while stochastic or plane-wave backgrounds and moving lenses generate distinctive, velocity-dependent deflection and delay terms in the weak field Kaiser and Jaffe (1997); Damour and Esposito-Farese (1998); Kopeikin and Schaefer (1999). Finally, oscillatory near-horizon structures can drive ring-scale astrometric modulations akin to ours, for example, superradiant bosonic clouds around spinning holes, which likewise predict QNM-frequency screen signatures in photon-ring observables, though sourced by matter fields rather than by the spacetime’s own even-parity QNM Chen et al. (2023).
What is still missing is a unified, purely analytical framework that projects directly onto the observer’s screen Johnson et al. (2020), is gauge-safe at first order, and continuously connects the far-field Born law to a time-dependent SDL with QNM-modulated coefficients, while demonstrating explicit matching between the two regimes. The present work fills this gap by deriving a single, line-of-sight Born kernel valid for all impact parameters and showing how the same kernel encodes the weak-field centroid wobble and the near-critical modulations, including the oscillation of the critical scale $b_{c}(t_{o})$, thereby providing a coherent bridge between ringdown spectroscopy and imaging observables. For the gravitational-wave analogue—namely, the image-type–dependent distortions of lensed GW waveforms even in the geometric-optics limit—see Dai and Venumadhav (2017).
We emphasize three theoretical payoffs of the unified treatment. First, weak-field centroid wobble, intermediate-field image multiplicity and delays, and strong-field photon-ring modulations are shown to be phase locked to the same QNM driver. Second, the link across regimes is explicit: the near-critical coefficients that govern the logarithmic law and the ring modulation are encoded in the near-critical expansion of the very amplitude that produces the weak-field $1/b$ correction. Third, all results follow from analytic asymptotics of a single line-of-sight kernel and therefore require no numerical ray tracing or data analysis.
@@ II.1 Schwarzschild + even-parity $(\ell,m)=(2,0)$ QNM
We take as background the Schwarzschild line element Chandrasekhar (1985)
$ds^{2}=-f(r)dt^{2}+f(r)^{-1}dr^{2}+r^{2}\left(d\theta^{2}+\sin^{2}\theta d\phi^{2}\right),$ (1)
@@
We consider a linear perturbation of amplitude $\varepsilon\ll 1$,
$g_{\mu\nu}(t,r,\theta,\phi)=g^{(0)}_{\mu\nu}(r)+\varepsilon\,h_{\mu\nu}(t,r,\theta,\phi),$ (3)
where $g^{(0)}_{\mu\nu}$ is the Schwarzschild metric in Eq. (1) and $h_{\mu\nu}$ is an even-parity, axisymmetric QNM with $(\ell,m)=(2,0)$. In the Regge-Wheeler-Zerilli (RWZ) gauge for $\ell\geq 2$, the nonvanishing components of an even-parity, axisymmetric perturbation can be written as Regge and Wheeler (1957); Zerilli (1970); Martel and Poisson (2005)
@@
$h_{\phi\phi}$ $=r^{2}K(t,r)\,P_{2}(\cos\theta)\,\sin^{2}\theta,$ (4)
where $H_{0},H_{1},H_{2},K$ are functions of $(t,r)$ and $P_{2}$ is the Legendre polynomial of degree $\ell=2$. Axisymmetry implies no $\phi$-dependence. Throughout the axisymmetric ($m=0$) analysis here, we use the Legendre basis $P_{\ell}(\cos\theta)$ for angular dependence. In particular, for $\ell=2$, $P_{2}(0)=-1/2$ for an equatorial observer. It is also convenient to package $(H_{0},H_{1},H_{2},K)$ into the gauge-invariant Zerilli master function $\Psi_{Z}(t,r)$, which fully characterizes the even-parity sector at first order.
Introducing the tortoise coordinate $r_{*}$ defined by Chandrasekhar (1985)
$\frac{dr_{*}}{dr}=f(r)^{-1},\qquad r_{*}=r+2M\ln\,\left(\frac{r}{2M}-1\right),$ (5)
@@
$\left(-\partial_{t}^{2}+\partial_{r_{*}}^{2}-V_{Z}(r)\right)\Psi_{Z}(t,r)=0,$ (6)
with the Zerilli potential Zerilli (1970); Moncrief (1974)
@@
$\lambda$ $\equiv\frac{(\ell-1)(\ell+2)}{2}.$ (7)
For the $(\ell,m)=(2,0)$ QNM, we impose homogeneous boundary conditions consistent with black-hole ringdown: purely ingoing at the event horizon and purely outgoing at null infinity. Writing $\Psi_{Z}(t,r)\propto e^{-i\omega t}\psi(r)$ with complex frequency $\omega=\omega_{R}+i\omega_{I}$ and $\omega_{I}<0$, the asymptotics are Berti et al. (2009)
@@
$\Psi_{Z}$ $\sim e^{-i\omega(t-r_{*})}\quad\text{for}\quad(r\to\infty).$ (8)
@@
At linear order in $\varepsilon$, the metric perturbation $h_{\mu\nu}$ can be reconstructed from $\Psi_{Z}$ through algebraic and first-derivative relations standard in the RWZ formalism; we defer the explicit formulas to Appendix A. For definiteness, we normalize $\Psi_{Z}$ so that the leading outgoing amplitude at large $r$ is unity, and we absorb the physical smallness of the perturbation into $\varepsilon$. With these conventions, Eq. (3) provides the unique axisymmetric, even-parity $\ell=2$ driving mode that will source the time dependence entering our unified lensing formalism in later sections.

### v2

# Time-Dependent Black Hole Lensing from Ringdown Quasinormal Mode

@@ Abstract
Is it possible to find imprints of a black hole ringdown through gravitational lensing? To address this question, we formulate an analytic description of weak-field and strong-deflection lensing of light in a time-dependent, perturbed Schwarzschild spacetime. The spacetime dynamics are modeled by a single, axisymmetric, even-parity quasinormal mode with $\ell=2$, $m=0$ and complex frequency $\omega$. Working to first order in a small perturbation amplitude while keeping background null geodesics exact, we derive a time-dependent line-of-sight (Born) expression for the screen-plane deflection measured by a static observer at large radius. From the same integral, an asymptotic expansion yields the familiar weak-field $1/b$ law with a ringdown-frequency correction that drives a harmonic centroid wobble, whereas a near-photon-sphere expansion produces a time-dependent generalization of the logarithmic strong-deflection limit with modulated coefficients, including a small oscillation of the critical impact parameter. An observer tetrad built from the background static frame ensures that all screen-plane quantities, such as centroid motion, multi-image hierarchy, and time delays, as well as photon-ring morphology, are gauge-safe at first order. We provide explicit matching across regimes, showing that the near-critical coefficients governing spacing and ring-radius modulations are encoded in the same Born kernel that controls the weak-field correction. This provides an analytic account of how ringdown-scale perturbations enter imaging observables, without resorting to numerical integration of null geodesics.
Gravitational lensing; black-hole ringdown; quasinormal modes; Zerilli perturbations; strong-deflection limit; centroid wobble
@@ I Introduction
Gravitational lensing by compact objects remains a powerful probe of strong-field gravity and the nature of dark compact sources [62, 23, 24, 17, 27, 41, 44, 5, 25, 31, 18, 42, 6, 58]. Foundational analyses of Schwarzschild lensing and its extensions to naked singularities established the basic phenomenology of multiple images, relativistic loops, and their observational signatures [47, 46]. Early work also highlighted how additional fields can imprint themselves on lensing observables [49] and connected these effects to broader questions such as Seifert’s conjecture [50]. The differential‐geometric underpinning of photon surfaces furnished a rigorous framework for locating unstable light rings that control both strong lensing and shadow formation [13]. Building on this, detailed predictions for relativistic images, time delays, and magnification centroids were developed and refined [51, 48]. Recent progress has quantified image distortions in Schwarzschild lensing and uncovered conservation properties of distortion measures, providing robust diagnostics that are less sensitive to astrophysical systematics [52, 54]. In parallel, precise relations between the photon-sphere radius and black-hole shadow have been extended to include cosmological constant corrections, clarifying how large-scale curvature perturbs near-horizon optics [2]. These theoretical advances feed directly into constraints on the compactness of supermassive dark objects at galactic centers, sharpening tests of black-hole paradigms with high-resolution imaging and timing data [53]. Realistic black hole lensing is rarely vacuum: ambient plasma and magnetic fields induce frequency-dependent deflection, reshape caustics, and shift photon rings and shadows [33, 38, 43]. Such chromatic and dispersive effects persist across rotating and regular spacetimes, complicating image interpretation [33, 38]. Beyond GR, braneworld geometries predict systematic deviations in strong-deflection and retrolensing observables, offering complementary tests [1].
@@
Gravitational lensing by black holes spans a wide dynamic range from weak deflections in the asymptotic region to the strong-field regime near the photon sphere [47, 35]. In parallel, the spacetime itself can be time-dependent, for instance, during black-hole ringdown when quasinormal modes (QNMs) drive metric perturbations with characteristic complex frequencies [8, 26]. These two features, wide spatial scales and intrinsic temporal modulation, are often modeled with separate tools that obscure their continuity. Our objective is to provide a single, purely analytical framework that treats both regimes within one description and that projects directly onto the observer’s screen in a gauge-safe manner.
@@
We concentrate on light propagation in the Schwarzschild geometry of mass $M$, perturbed at first order by a single axisymmetric even-parity Zerilli mode with angular numbers $\ell=2$, $m=0$ and complex frequency $\omega=\omega_{R}+i\omega_{I}$ with $\omega_{I}<0$ [61, 32]. For the lowest Schwarzschild overtones, the ringdown scale is set by $M^{-1}$ (see standard QNM tables in Ref. [8]). This choice provides the simplest dynamical driver of ringdown that preserves axial symmetry, minimizes technical overhead, and still captures the essential physics of a decaying spacetime oscillation. Throughout, we work perturbatively in a small, dimensionless amplitude $\varepsilon$ and keep the unperturbed null geodesics exact, so that all time dependence arises from the interaction of the QNM with the photon’s background trajectory [27, 29].
@@
Our central construct is a time-dependent line-of-sight map that gives the deflection measured by a static observer at a large areal radius. The map is evaluated along the unperturbed null geodesic labeled by its impact parameter $b$. In the far field, it reproduces the familiar weak-bending law ($1/b$) [59, 60] with an $\varepsilon$-suppressed, frequency-resolved correction that produces a harmonic motion of the image centroid on the screen. In the near-critical regime $b\to b_{c}^{+}$, with $b_{c}$ the Schwarzschild critical impact parameter, the same map yields a time-dependent generalization of the logarithmic strong-deflection relation [9, 41]. The coefficients of this relation become slowly varying functions of the observer’s time, all oscillating at the QNM frequency. The critical scale itself acquires a small modulation that controls the most singular response near the photon ring [11, 23, 22, 3, 4].
@@
A key structural feature of our approach is the use of an observer tetrad fixed by the background static frame [32, 23]. Screen angles are constructed directly from locally measured photon momenta. At first order in $\varepsilon$, this guarantees that centroid shifts, image positions, and ring morphology are insensitive to gauge transformations that would otherwise act on the metric perturbation. In this way, the imaging diagnostics we develop are bona fide observables at the order of interest.
@@
Recent work has begun to probe time-dependent gravitational lensing during black-hole relaxation. In particular, Zhong, Cardoso, and Chen [62] have used backward ray tracing in a relaxing Kerr background to show that the photon deflection inherits the characteristic ringdown pattern at intermediate times and exhibits a $t_{o}^{-3}$ late-time decay, thereby establishing the qualitative imprint of QNM dynamics on light propagation [36]. Earlier analyses of time-dependent deflection by gravitational waves or localized, retarded sources also demonstrated oscillatory signatures in lensing and timing, albeit in regimes far from the photon sphere and without a black-hole ringdown setting [24, 17, 57, 34]. At the same time, the strong-deflection limit (SDL) for stationary spacetimes is well understood, including the logarithmic divergence and its coefficients ($\bar{a},\bar{b},b_{c}$) that control image hierarchies and photon-ring morphology [9, 41]. Classic studies of light propagation in nonstationary fields also delineate key limits: in particular, for localized sources, the wave-zone and intermediate-zone contributions cancel so that the leading time-dependent deflection is near-zone and falls as $b^{-3}$ (contrasting with the $1/b$ Born imprint we obtain from a global QNM on a black-hole background), while stochastic or plane-wave backgrounds and moving lenses generate distinctive, velocity-dependent deflection and delay terms in the weak field [24, 17, 27]. Finally, oscillatory near-horizon structures can drive ring-scale astrometric modulations akin to ours, for example, superradiant bosonic clouds around spinning holes, which likewise predict QNM-frequency screen signatures in photon-ring observables, though sourced by matter fields rather than by the spacetime’s own even-parity QNM [12].
A central goal of this paper is to express screen-plane lensing observables for a time-dependent perturbation in a form that applies from the weak-field regime to the strong-deflection limit. We do so by deriving a single line-of-sight Born kernel that is valid for all impact parameters at $\mathcal{O}(\varepsilon)$, and by extracting both the far-field $1/b$ correction and a time-dependent extension of the SDL through controlled asymptotic expansions. The same kernel also governs the modulation of the critical scale $b_{c}(t_{o})$ and the associated near-critical imaging diagnostics. For the gravitational-wave analogue, namely, the image-type-dependent distortions of lensed GW waveforms even in the geometric-optics limit, see Ref. [16]. SUch a treatment allows us to track how a single QNM perturbation induces a shared phase in weak-field centroid motion and in near-critical photon-ring modulations, with both regimes governed by analytic asymptotics of the same line-of-sight kernel.
@@ II.1 Schwarzschild + even-parity $(\ell,m)=(2,0)$ QNM
We take as background the Schwarzschild line element [40, 56, 11]
$ds^{2}=-f(r)dt^{2}+f(r)^{-1}dr^{2}+r^{2}\left(d\theta^{2}+\sin^{2}\theta d\phi^{2}\right),$ (II.1)
@@
We consider a linear perturbation of amplitude $\varepsilon\ll 1$ [37, 30],
$g_{\mu\nu}(t,r,\theta,\phi)=g^{(0)}_{\mu\nu}(r)+\varepsilon\,h_{\mu\nu}(t,r,\theta,\phi),$ (II.3)
where $g^{(0)}_{\mu\nu}$ is the Schwarzschild metric in Eq. (II.1) and $h_{\mu\nu}$ is an even-parity, axisymmetric QNM with $(\ell,m)=(2,0)$. In the Regge-Wheeler-Zerilli (RWZ) gauge for $\ell\geq 2$, the nonvanishing components of an even-parity, axisymmetric perturbation can be written as [37, 61, 32, 30]
@@
$h_{\phi\phi}$ $=r^{2}K(t,r)\,P_{2}(\cos\theta)\,\sin^{2}\theta,$ (II.4)
where $H_{0},H_{1},H_{2},K$ are functions of $(t,r)$ and $P_{2}$ is the Legendre polynomial of degree $\ell=2$. Axisymmetry implies no $\phi$-dependence. Throughout the axisymmetric ($m=0$) analysis here, we use the Legendre basis $P_{\ell}(\cos\theta)$ for angular dependence. In particular, for $\ell=2$, $P_{2}(0)=-1/2$ for an equatorial observer. It is also convenient to package $(H_{0},H_{1},H_{2},K)$ into the gauge-invariant Zerilli master function $\Psi_{Z}(t,r)$, which encodes the dynamical content of the even-parity sector at first order.
Introducing the tortoise coordinate $r_{*}$ defined by [11]
$\frac{dr_{*}}{dr}=f(r)^{-1},\qquad r_{*}=r+2M\ln\,\left(\frac{r}{2M}-1\right),$ (II.5)
@@
$\left(-\partial_{t}^{2}+\partial_{r_{*}}^{2}-V_{Z}(r)\right)\Psi_{Z}(t,r)=0,$ (II.6)
with the Zerilli potential [61, 32]
@@
$\lambda$ $\equiv\frac{(\ell-1)(\ell+2)}{2}.$ (II.7)
For the $(\ell,m)=(2,0)$ QNM, we impose homogeneous boundary conditions consistent with black-hole ringdown: purely ingoing at the event horizon and purely outgoing at null infinity. Writing $\Psi_{Z}(t,r)\propto e^{-i\omega t}\psi(r)$ with complex frequency $\omega=\omega_{R}+i\omega_{I}$ and $\omega_{I}<0$, the asymptotics are [26, 8, 55, 28]
@@
$\Psi_{Z}$ $\sim e^{-i\omega(t-r_{*})}\quad\text{for}\quad(r\to\infty).$ (II.8)
@@
At linear order in $\varepsilon$, the metric perturbation $h_{\mu\nu}$ can be reconstructed from $\Psi_{Z}$ through algebraic and first-derivative relations standard in the RWZ formalism; we defer the explicit formulas to Appendix A. For definiteness, we normalize $\Psi_{Z}$ so that the leading outgoing amplitude at large $r$ is unity, and we absorb the physical smallness of the perturbation into $\varepsilon$. With these conventions, Eq. (II.3) provides a definite axisymmetric, even-parity $\ell=2$ driving mode that sources the time dependence used in our lensing construction in later sections.